### v1

# The Exact Uncertainty Relation and Geometric Speed Limits in Krylov Space

Abstract
We show that Hall’s exact uncertainty relation acquires a simple geometric form in the Krylov basis generated by the Liouvillian. In this canonical operator frame, the uncertainty equality implies that the operator amplitude vector evolves on the unit Krylov sphere with constant speed fixed solely by the first Lanczos coefficient. This yields an exact linear bound on geometric operator evolution, independent of higher Lanczos coefficients and valid for arbitrary Hamiltonians, integrable or chaotic. Our results provide the first unified geometric interpretation of exact quantum speed limits and operator growth, identifying the first Lanczos coefficient as the intrinsic speed scale of quantum dynamics.
This work is dedicated to Sudipta Mukherji on the occasion of his 60th birthday.
I Introduction
The uncertainty principle is a defining feature of quantum mechanics, limiting the precision with which certain pairs of observables can be simultaneously measured. Beyond the canonical position–momentum relation, energy–time uncertainty relations Mandelstam and Tamm (1945); Margolus and Levitin (1998) imply fundamental limits on how rapidly a quantum state may evolve, known as quantum speed limits (QSLs). Such limits play a central role in quantum control, quantum information processing, and quantum computation.
Most uncertainty relations take the form of inequalities and therefore yield lower bounds on the time required for a given evolution. In contrast, Hall derived a state-dependent exact uncertainty relation Hall (2001) for two observables $A$ and $B$ in a pure state $\rho=|\psi\rangle\langle\psi|$,
$\delta_{B}A\,\Delta B_{nc}=\frac{1}{2},$ (1)
where $\Delta B_{nc}$ denotes the nonclassical component of $B$, and $\delta_{B}A$ is a Fisher-information-like functional of $A$. Although explicitly state dependent, this relation yields an equality-type QSL once the nonclassical generator is identified in a suitable basis. Indeed, Ref. Pati et al. (2023) used Hall’s relation to construct an exact QSL for quantum dynamics.
In this work, we show that operator dynamics admits an exact geometric formulation in Krylov space, in which quantum speed limits arise as kinematic identities. Specifically, we demonstrate that Hall’s exact uncertainty relation acquires a particularly simple and universal form in the Krylov basis, where the nonclassical part of the Liouvillian coincides with the full generator. As an immediate consequence, Lanczos representation of the Liouvillian Lanczos (1950); Viswanath and Müller (2008) reveals operator evolution on the unit Krylov sphere with a constant speed fixed solely by the first Lanczos coefficient. This could be identified as a geometric Lieb–Robinson–type bound intrinsic to Krylov dynamics which remains exact even when the Lanczos coefficients grow without bound, as in typical chaotic systems. In our work, we demonstrate this by introducing the motion of amplitude fronts in Krylov index space which is distinct from the geometric motion of the operator state on the Krylov sphere. We show that even when the operator front accelerate rapidly in Krylov index space, the actual geometric motion of the operator state remains strictly bounded, thereby resolving the apparent tension between fast Krylov growth and finite causal structure.
Recent work has shown that Krylov space provides a powerful framework for characterizing operator growth and information spreading in many-body quantum systems Roberts et al. (2018); Barbón et al. (2019). Our results place these developments in a concrete geometric setting and identify a basis-independent speed law that governs the operator dynamics. The novelty of this speed law lies in the fact that its universality remains valid across the choice of a Hamiltonian, be it chaotic or integrable. It is interesting to note that while the higher Lanczos coefficients do not affect the speed, they do govern the geometry of the trajectory—its curvature and departure from geodesic motion, leading to universal constraints such as bounds on return amplitudes.
This paper is organized as follows. Section 1 reviews the Krylov construction, derives the constant-speed identity, and shows its equivalence to Hall’s relation in Krylov space. Section 2 develops the geometric picture: curvature and torsion, geodesics, the Krylov light cone, tail bounds, and quadratic invariants. Section 3 contains our conclusions.
II Krylov Bases as Optimal Frames and the Exact Speed Law
[…]
In this section, we review the operator Krylov construction  [^1] In what follows we will consider operator growth, though it can be generalized to state spreading Balasubramanian et al. (2022); Alishahiha and Banerjee (2022); Caputa et al. (2024); Alishahiha (2023) (See Alishahiha and Banerjee ). and show that Krylov space is precisely the basis in which the exact uncertainty relation of Hall is saturated. This directly leads to an exact speed law, showing that the “velocity” of operator evolution in Krylov space is time-independent and given by the first Lanczos coefficient $b_{1}$.
Consider an operator ${\cal O}(0)$ evolving in the Heisenberg picture,
${\cal O}(t)=e^{-iHt}\,{\cal O}(0)\,e^{iHt}.$ (2)
Although ${\cal O}(t)$ is an element of the full operator algebra, its time evolution is restricted to the Krylov subspace
${\cal B}_{\mathcal{O}}(H)=\{{\cal O},[H,{\cal O}],[H,[H,{\cal O}]],\dots\},$ (3)
whose dimension is typically much smaller than $d^{2}$.
Introducing the Liouvillian superoperator ${\cal L}{\cal O}=[H,{\cal O}]$, the Heisenberg evolution can be compactly written as ${\cal O}(t)=e^{-it{\cal L}}{\cal O}(0)$. which solves the operator Schrödinger equation.
To construct a convenient basis for ${\cal B}_{\mathcal{O}}(H)$, we introduce the operator–state mapping $|{\cal O})$, with inner product $(A|B)=\mathrm{Tr}(A^{\dagger}B)$, and apply Gram–Schmidt orthogonalization to the non-orthogonal set $\{{\cal L}^{n}{\cal O}\}$. The resulting orthonormal vectors $\{|K_{n}\rangle\}_{n=0}^{D_{\mathcal{O}}-1}$ with $D_{\mathcal{O}}\leq d^{2}-d+1$, satisfy the Lanczos recursion Viswanath and Müller (2008)
$|K_{0}\rangle=\frac{|\mathcal{O}(0))}{||\mathcal{O}(0))|},\quad{\cal L}|K_{n}\rangle=b_{n+1}|K_{n+1}\rangle+b_{n}|K_{n-1}\rangle,$ (4)
which is tridiagonal and minimal in bandwidth. This bandwidth-minimization characterizes Krylov space as an optimal basis for describing operator growth. Here $b_{0}=0$ and $b_{n\geq 1}\geq 0$ are the Lanczos coefficients that characterize the rate of operator spreading.
Let us expand the time-evolved operator in this basis,
$|{\cal O}(t))=\sum_{n=0}^{D_{\mathcal{O}}-1}i^{n}\phi_{n}(t)\,|K_{n}\rangle,\;\;\;\;\sum_{n=0}^{D_{\cal O}-1}\phi_{n}^{2}(t)=1.$ (5)
Indeed, this implies that the Heisenberg evolution of the operator $|{\cal O}(t))$ can be viewed as a point moving along a trajectory on the unit sphere in Krylov space, dubbed Krylov sphere, with coordinates $\{\phi_{n}(t),\;n=0,1,\dots,\,D_{\cal O}-1\}$ satisfying
$\dot{\phi}_{n}(t)=b_{n}\phi_{n-1}(t)-b_{n+1}\phi_{n+1}(t),$ (6)
with boundary conditions $b_{0}=0$ and $\phi_{-1}=0$.
Taking the square of the operator Schrödinger equation $i\dot{\cal O}(t)={\cal L}{\cal O}(t)$, yields the fundamental identity
$\sum_{n=0}^{D_{\mathcal{O}}-1}\dot{\phi}_{n}^{2}(t)=\left({\mathcal{O}(t)}|{\cal L}^{2}|{\mathcal{O}(t)}\right)=b_{1}^{2},$ (7)
valid for all times $t$. This immediately sets an intrinsic Krylov velocity
$v_{K}(t)=\sqrt{\sum_{n}\dot{\phi}_{n}^{2}(t)},$ (8)
which is strictly constant in time along the trajectory on the sphere (7) and equal to the first Lanczos coefficient $b_{1}$. In Sec. III2, we demonstrate the roles of higher Lanczos coefficients $b_{n\geq 2}$ in determining the curvature and torsion of the trajectory in Krylov space, thus providing a complete description of Krylov geometry.
Let us conclude the section by demonstrating the equivalence between (7) and Hall’s exact uncertainty relation (1). For this we decompose the Liouvillian into classical and nonclassical components relative to $\varrho(t)=|{\cal O}(t))({\cal O}(t)|$. It turns out that the classical part of the Liouvillian vanishes identically (for details see supplemental material) so that its non-classical part ${\cal L}_{nc}={\cal L}$. Using (7) we get
$\Delta{\cal L}_{nc}=\sqrt{\langle{\cal L}^{2}\rangle-\langle{\cal L}\rangle^{2}}=b_{1}.$ (9)
[…]
$(\delta_{{\cal L}}A)^{-2}=\sum_{n=0}^{D_{\cal O}-1}\frac{\langle K_{n}|i\,[{\cal L},\varrho]|K_{n}\rangle^{2}}{\langle K_{n}|\varrho|K_{n}\rangle}=4\sum_{n=0}^{D_{\cal O}-1}\dot{\phi}_{n}^{2}(t).$ (10)
Substituting (9) and (10) into the exact uncertainty relation (1), we recover precisely (7). This shows that the exact uncertainty relation becomes identity (7) in the Krylov basis. Intermediate steps of this derivation can be found in the supplemental material.
III Geometry and Dynamics on the Krylov Sphere
III.1 Geometry of Krylov Sphere
Let us consider the motion along a trajectory $\Phi(t)=\{\phi_{0}(t),\phi_{1}(t),\dots\ ,\phi_{D_{\cal O}-1}(t)\}$ on the Krylov sphere. Equation (7) ensures that the speed associated with this motion is constant,
$|\dot{\Phi}(t)|^{2}=\sum_{n=0}^{D_{\cal O}-1}\dot{\phi}_{n}^{2}(t)=b_{1}^{2},$ (11)
so that the total arc–length traversed between times $t_{0}$ and $t_{1}$ is
$\ell_{K}(t_{0},t_{1})=\int_{t_{0}}^{t_{1}}dt\,|\dot{\Phi}(t)|=b_{1}(t_{1}-t_{0}).$ (12)
Differentiating (11) yields $\dot{\Phi}\cdot\ddot{\Phi}=0$, so the acceleration is everywhere orthogonal to the velocity. Hence the Krylov trajectory is a curve of constant speed and nonzero curvature on the sphere—very similar to uniform circular motion, though not necessarily periodic. This geometric picture is reminiscent of the ideas presented in Brown and Susskind (2019).
Because the Krylov trajectory $\Phi(t)$ moves at constant speed on the unit sphere, standard Frenet–Serret constructions apply directly, allowing us to define curvature and torsion in terms of the first few time derivatives of $\Phi(t)$. The curvature of the Krylov trajectory is given by
$\kappa(t)=\frac{|\dot{\Phi}\times\ddot{\Phi}|}{|\dot{\Phi}|^{3}}=\sqrt{1+\frac{b_{2}^{2}}{b_{1}^{2}}}\,$, (13)
where we have also used that fact that $\dot{\Phi}\cdot\ddot{\Phi}=0$ and $|\ddot{\Phi}|^{2}=b_{1}^{2}(b_{1}^{2}+b_{2}^{2})$. One may also define the torsion using the triple of derivatives  [^2] Practically, the determinant can be computed by forming the $3\times 3$ Gram matrix $G_{ij}=v_{i}\!\cdot\!v_{j}$ with $v_{1}=\dot{\Phi},v_{2}=\ddot{\Phi},v_{3}=\dddot{\Phi}$ and then using $\det[\dot{\phi},\ddot{\phi},\dddot{\phi}]^{2}=\det G$.
$\tau(t)=\frac{\det[\dot{\Phi},\ddot{\Phi},\dddot{\Phi}\big]}{|\dot{\Phi}\times\ddot{\Phi}|^{2}}=\frac{b_{2}b_{3}}{b_{1}^{2}\sqrt{b_{1}^{2}+b_{2}^{2}}}\,.\,.$ (14)
It is then clear that if the dynamics never populates a third independent Krylov direction (e.g. at most $b_{1},b_{2}\neq 0$ while $b_{n\geq 3}=0$), then $\tau=0$. In general, when higher Lanczos coefficients are nonzero, the operator trajectory explores a higher-dimensional subspace of the Krylov sphere and acquires nonzero curvature and torsion. These quantities characterize the bending and twisting of the operator’s path on the Krylov sphere and are dynamical invariants determined by the Lanczos coefficients.
Using (11) it is easy to check that The operator’s path is a geodesic iff $\ddot{\Phi}=-b_{1}^{2}\Phi$ which in terms of components read $\ddot{\phi}_{n}=-b_{1}^{2}\phi_{n}$, whose left hand side can be directly computed by differentiating and successively using (6)
$\ddot{\phi}_{n}=-\big(b_{n}^{2}+b_{n+1}^{2}\big)\phi_{n}+b_{n}b_{n-1}\,\phi_{n-2}+b_{n+1}b_{n+2}\,\phi_{n+2}\,$, (15)
from which we immediately conclude that Lanczos dynamics is a geodesic motion on the Krylov sphere when the coefficients for all $n$ satisfy
$b_{n}b_{n-1}\,\phi_{n-2}+b_{n+1}b_{n+2}\,\phi_{n+2}=0\,,\;\;\;b_{n}^{2}+b_{n+1}^{2}=b_{1}^{2}\,.$ (16)
For generic nonzero $\Phi$ the only consistent and physically relevant solution of (16) is the truncation $b_{1}\neq 0$, and $b_{n\geq 2}=0$, i.e. dynamics confined to a two-dimensional Krylov subspace with $\kappa=1$. Thus the trajectory is geodesic if and only if the Krylov chain truncates to two dimensions. Whenever higher Lanczos coefficients are nonzero, the motion is intrinsically non-geodesic and acquires a non-trivial curvature and torsion  [^3] This may be related to the observation that the Krylov complexity cannot be considered as a distance Aguilar-Gutierrez and Rolph (2024). Growth of $\kappa$ thus signals departure from geodesicity and increased operator spreading into higher Krylov layers. In the supplemental material we discuss simple examples of Hamiltonians which give rise to geodesic and non-geodesic trajectories on the Krylov sphere.
An immediate consequence of the non-geodesicity is a universal bound on the return amplitude $\phi_{0}(t)=\left(K_{0}|{\mathcal{O}}(t)\right)$. Let $\theta(t)$ be the geodesic distance between $\Phi(t)$ and $\Phi(0)$, then $\theta(t)=\arccos(\phi_{0}(t))$. Since geodesic distance can never exceed arc–length, from (12) we have $\theta(t)\leq b_{1}t$, resulting in an exact lower bound on the return amplitude  [^4] Using the fact that the first Lanczos coefficient is given by the energy variance (9), the lower bound on the return amplitude may be interpreted as the Mandelstam–Tamm quantum speed limit bound Mandelstam and Tamm (1945) $\frac{\theta(t)}{\Delta\mathcal{L}_{\rm nc}}\leq t$.
$\phi_{0}(t)\;\geq\;\cos(b_{1}t),\;\qquad 0\;\leq b_{1}t\;\leq\pi/2.$ (17)
III.2 Light cone in the Krylov sphere
Equation (12) provides a strict geometric constraint on the evolution of the operator state on the Krylov sphere. Because the Lanczos recursion is strictly nearest–neighbor, the spreading of amplitude in Krylov space is sequential: level $n$ can only be reached after propagating through levels $0,1,\dots,n-1$. It is therefore natural to expect that the geometric bound on arc–length implies a light–cone–like constraint on the growth of the Krylov index. In what follows, we make this emergent Krylov locality precise by introducing a time–dependent front $n(t)$, defined as the largest Krylov level with appreciable amplitude. Using the operator norm bound on the Liouvillian that follows directly from its tridiagonal structure, we derive a rigorous bound on $n(t)$ of the form (see the supplemental material for details of the derivation)
$|\phi_{n}(t)|\leq e^{-n\log\frac{n}{v_{\mathrm{op}}t}+n+v_{\mathrm{op}}t},$ (18)
where $v_{\mathrm{op}}\equiv\max_{n}(b_{n}+b_{n+1})$. This results in an exponential decay when $n\gtrsim c\,v_{\mathrm{op}}t$ with $c\gtrsim 3.6$. Thus one obtains a Lieb–Robinson–type light cone in Krylov index space,
$\phi_{n}(t)\approx 0,\qquad n>{v_{\mathrm{op}}}t,$ (19)
upto a numerical constant. As an illustrative example, consider the case in which all Lanczos coefficients are constant, $b_{n}=b$ for all $n$. In this case, the recursion relation (6) can be solved exactly in terms of Bessel functions, yielding
$\phi_{n}(t)=(-1)^{n}J_{n}(2bt).$ (20)
The resulting wavefunction is peaked around $n\approx 2bt$ and decays exponentially outside this region, demonstrating that the Krylov light–cone bound in index space is essentially sharp.
A subtle issue appears in infinite-dimensional settings where the Lanczos coefficients diverge, $b_{n}\to\infty$ as $n\to\infty$, so that formally
$\max_{n}(b_{n}+b_{n+1})=\infty.$ (21)
This indeed occurs, for example, in the harmonic oscillator where $b_{n}\sim\sqrt{n}$ and in chaotic systems, where one typically finds $b_{n}\sim\alpha n$. At first sight, this seems to be in direct tension with the exact geometric speed law $\ell_{K}(t)=b_{1}t$, which strictly limits the distance traversed by the operator state on the Krylov sphere.
The resolution of this apparent paradox is conceptual rather than technical. It arises from distinguishing two fundamentally different notions of motion: (i) the geometric motion of the operator state on the Krylov sphere, and (ii) the motion of the amplitude “front” in Krylov index space  [^5] The notion of a propagating wavefront in Krylov index space and its accelerated motion for growing Lanczos coefficients has appeared previously in studies of Krylov complexity and operator growth, see e.g. Refs. Rabinovici et al. (2023); Ambrosini et al. (2025). Our emphasis here is different: we distinguish this index-space redistribution from the true geometric motion of the operator state, which we show is exactly bounded by a linear arc-length law controlled solely by the first Lanczos coefficient. . The measure of (i) is the true physical displacement of the operator along a unit-speed trajectory on the Krylov sphere and is governed by the arc–length $\ell_{K}(t)$. This motion therefore defines a natural Lieb–Robinson–type velocity in Krylov space, $v_{\mathrm{op}}=b_{1}$. By contrast, (ii) merely tracks how the wavefunction is distributed among basis elements labeled by the discrete index $n$. Clearly, these two notions need not scale in the same way. In particular, when the Lanczos coefficients grow with $n$, successive Krylov levels correspond to progressively smaller angular displacements on the sphere. As a result, the operator may traverse many Krylov levels while covering only a small geometric distance. Thus, even when the front $n(t)$ accelerates, the geometric motion remains strictly bounded by $\ell_{K}(t)=b_{1}t$, and no causal constraint is violated.
In order to demonstrate this, let us note once again that the Lanczos recursion couples only nearest neighbors, and therefore the characteristic time to transfer amplitude from level $n$ to $n+1$ can be estimated as $\Delta t_{n}\sim 1/{b_{n}}$. This immediately implies the minimal-time constraint for reaching level n
$\sum_{m=1}^{n(t)}\frac{1}{b_{m}}\;\leq\;t,$ (22)
or, in the continuum approximation QSL ,
$\int_{1}^{n(t)}\frac{dn}{b_{n}}\;\lesssim\;t\,$, (23)
which gives the maximal Krylov index reachable by time $t$, independent of interference or combinatorial enhancements.
Geometrically, the angular step at level $n$ scales as $\theta_{n}=b_{1}\Delta t_{n}\sim b_{1}/b_{n}$, so Eq. (22) is equivalent to
$\sum_{m=1}^{n(t)}\theta_{m}\leq b_{1}t.$ (24)
For bounded $b_{n}$ this yields linear growth $n(t)\leq b_{1}t$, while unbounded $b_{n}$ leads to nonlinear fronts, e.g. (Further details are given in Supplementary materials),
$b_{n}\sim\alpha n\Rightarrow n(t)\lesssim e^{\alpha t},\qquad b_{n}\sim\alpha\sqrt{n}\Rightarrow n(t)\lesssim\tfrac{(\alpha t)^{2}}{4}.$ (25)
Thus, even when Lanczos coefficients diverge and the front accelerates, the operator’s geometric displacement remains strictly bounded by $\ell_{K}(t)=b_{1}t$. The true light cone is linear and controlled solely by $b_{1}$, resolving the apparent conflict between unbounded Lanczos growth and finite causal structure.
III.3 Invariants of the motion
The Krylov dynamics (6) can be compactly reexpressed as
$\dot{\Phi}=A\Phi,\qquad\Phi(t)=e^{tA}\Phi_{0},$ (26)
where the Krylov hopping matrix $A$ is a real skew-symmetric tridiagonal matrix with entries $A_{n,n-1}=b_{n},\;A_{n,n+1}=-b_{n+1}$, and all other $A_{nm}=0$.
Let us define the scalar functional
$\mathcal{I}(t)=\Phi^{T}I\Phi,$ (27)
which is conserved whenever $[A,I]=0$. Trivial examples include functions of $A$, which in the operator language yield identities such as $(\mathcal{O}(t)|\mathcal{L}^{2n}|\mathcal{O}(t))={\mathrm{constant}}$ (and vanishing for odd powers). The latter follows from the fact that, in our notation, the Krylov hopping matrix $A$ is related to the Liouvillian matrix $L$ by a similarity transformation
$A=-i\,S^{-1}LS,\qquad S=\mathrm{diag}(1,i,i^{2},i^{3},\dots).$ (28)
where $L$ is the symmetric Lanczos representation of the Liouvillian $\mathcal{L}$ in the Krylov basis.
More generally, using the canonical form of real skew-symmetric matrices, one finds that the commuting symmetric matrices $I$ are precisely those that are block-diagonal with each $2\times 2$ rotation block proportional to the identity. Thus there are infinitely many quadratic constants of motion in an infinite-dimensional Krylov space.
Thus the Krylov dynamics is a linear orthogonal flow with many quadratic constants of motion and hence is integrable in the linear algebraic sense. This does not contradict, but rather underpins, the rich and often rapid growth of nonlinear complexity measures: those are nonlinear functionals of $\Phi$ and so can increase even when the underlying linear flow is nonchaotic. Understanding operator growth therefore requires combining this linear geometric picture with the Lanczos data $\{b_{n}\}$ and the nonlinear nature of the observables of interest.
Although the Krylov evolution is always integrable, this does not imply that the underlying quantum Hamiltonian is integrable. This structure may offer a new approach to diagnosing quantum chaos.
For $I_{nm}=n\delta_{nm}$ the scalar $\mathcal{I}$ reduces to Krylov complexity, $\mathcal{C}(t)=\sum_{n}n\,\phi_{n}^{2}(t)$. our geometric construction yields a remarkably simple and universal bound on Krylov complexity:
$\mathcal{C}(t)=\sum_{n}n\,\phi_{n}^{2}(t)\;\leq\;n(t).$ (29)
By contrast, the bound in Ref. Hörnedal et al. (2022) constrains the growth rate $\bigl|\partial_{t}\mathcal{C}(t)\bigr|$ via an uncertainty-type relation involving the instantaneous spread $\Delta\mathcal{C}(t)$, namely $\bigl|\partial_{t}\mathcal{C}(t)\bigr|\leq 2b_{1}\,\Delta\mathcal{C}(t)$. That bound is state- and fluctuation-dependent, becoming weak when $\Delta\mathcal{C}(t)$ grows large.
Hence our result has two pronounced advantages. First, it is state-independent and gives an absolute, time-integrated upper bound on complexity for all times. Second, it avoids any dependence on higher moments or variances of the Krylov distribution; instead it relies solely on the first Lanczos coefficient $b_{1}$.
IV Conclusion
We have shown that operator dynamics admits a natural geometric description in Krylov space, in which quantum speed limits, operator growth, and causal bounds can be understood in a unified and exact manner. In the Krylov basis, the operator state evolves on the unit Krylov sphere with a constant geometric speed set by the first Lanczos coefficient $b_{1}$. This implies linear growth of the accumulated arc length, $\ell_{K}(t)=b_{1}t$, and establishes an exact Lieb–Robinson–type bound on operator propagation in Krylov space, which is independent of microscopic locality and valid for all Hamiltonians.
A central result of this work is the resolution of the apparent tension between this linear light cone and the rapid growth of operators observed in chaotic systems. When the Lanczos coefficients increase with index, the operator front—the Krylov level carrying maximal weight—may accelerate and reach very large values of $n(t)$. However, this accelerated front does not correspond to superluminal propagation: each step in Krylov index costs a finite geometric distance, and large Lanczos coefficients only become dynamically relevant once the operator reaches the corresponding level. As a result, even in systems with unbounded $b_{n}$, the causal structure remains strictly controlled by $b_{1}$.
While $b_{1}$ fixes the speed, the higher Lanczos coefficients $b_{n\geq 2}$ control the geometry of the trajectory. If the Lanczos chain truncates at $n=1$, the motion is geodesic and reduces to uniform great-circle rotation, reproducing two-level Rabi dynamics. In generic systems, nonzero $b_{n\geq 2}$ induce non-geodesic motion and operator spreading, leading to universal constraints such as the bound on the return amplitude $\phi_{0}(t)\geq\cos(b_{1}t)$.
Finally, we clarified the relation between chaos and integrability in Krylov dynamics. The tridiagonal skew-symmetric structure of the Liouvillian guarantees integrability in a linear-algebraic sense and enforces strong geometric constraints even for chaotic Hamiltonians. At the same time, nonlinear measures of operator complexity may grow rapidly, demonstrating how accelerated front propagation and chaotic operator growth can coexist with an exact geometric light cone.
Together, these results identify the Krylov basis as an intrinsic geometric frame in which operator growth, front acceleration, and causal bounds become transparent, with the first Lanczos coefficient emerging as the fundamental velocity scale governing operator dynamics.
Acknowledgements
We would like to thank Mohammad Reza Tanhayi, Mohammad Javad Vasli for many insightful discussions on various aspects of complexity. We would like to thank Arun K. Pati for his comments on an earlier version of this manuscript. The work of M. A. is supported by the Iran National Science Foundation (INSF) under Project No. 4023620.
Supplemental Material for: The Exact Uncertainty Relation and Geometric Speed Limits in Krylov Space
Appendix A Hall uncertainty
In fact, for any two operators $A$ and $B$ representing two observables in a given quantum system described by a pure state $\rho=|\psi\rangle\langle\psi|$, the corresponding exact uncertainty relation may be written as follows (setting $\hbar=1$)
$\delta_{B}A\Delta B_{nc}=\frac{1}{2},$ (S1)
where
$(\delta_{B}A)^{-2}=\sum_{k}\frac{\langle a_{k}|\frac{i}{\hbar}[B,\rho]|a_{k}\rangle^{2}}{\langle a_{k}|\rho|a_{k}\rangle},$ (S2)
and $B_{nc}$ is non-classical part of the operator $B$ defined by $B_{nc}=B-B_{cl}$ where the classical part is given by
$B_{cl}=\sum_{k}\frac{\langle a_{k}|\frac{1}{2}\{B,\rho\}|a_{k}\rangle}{\langle a_{k}|\rho|a_{k}\rangle}\;|a_{k}\rangle\langle a_{k}|\,.$ (S3)
Here $\{|a_{k}\rangle\}$ is the eigenstates of the operator $A$.
The idea behind this proposal is the fact that for a given operator only the part which is non-classical gives a contribution to an uncertainty relation. In fact for a pair of non-commuting operator $\{A,B\}$ the classical part $B_{cl}$ is the best estimate one may make for observable $B$ for a given state that is compatible with the measurement of observable $A$. In other words the classical part of $B$ commutes with $A$. Therefore quantum effects are encodded in the non-classical part of $B$. It is also easy to see that Hall (2001) $\langle B\rangle=\langle B_{cl}\rangle,\langle B_{nc}\rangle=0$ and
$\langle B^{2}\rangle=\langle B_{cl}^{2}\rangle+\langle B^{2}_{nc}\rangle\,.$ (S4)
It is worth noting that the price we paid to get an equality is that the uncertainty relation (1) is state dependent which may not be so restrictive, as when it comes to an actual experiment we always start with a prepared quantum state. Moreover, for non-pure states the above uncertainty relation also reduces to an inequality Hall (2001).
Based on this uncertainty relation, recently and exact quantum speed limit has been proposed in Pati et al. (2023). For a $d$-dimensional quantum system, the time required to transport a given state $|\psi(0)\rangle$ to the target state $|\psi(T)\rangle$ via a unitary dynamics generated by the Hamiltonian $H$ is given by Pati et al. (2023)
$T=\frac{\hbar\ell(T)}{\langle\Delta H_{nc}\rangle_{T}}\,.$ (S5)
Here the length $\ell$ is given by
$\ell(T)=\int_{0}^{T}dt\,\sqrt{\sum_{k=0}^{d-1}\left(\frac{d|c_{k}|}{dt}\right)^{2}}$ (S6)
with $c_{i}$ being the expansion coefficients of $|\psi(t)\rangle=\sum_{k}c_{k}|a_{k}\rangle$. The non-classical part of the Hamiltonian is also computed in this basis. $\langle\Delta H_{nc}\rangle_{T}$ is time average of the non-classical part which is time-dependent even though the full Hamiltonian may not be time-dependent.
Following Hall (2001) decompose the Liouvillian into classical and nonclassical components relative to
$\varrho(t)=|{\cal O}(t))({\cal O}(t)|.$ (S7)
Using
$\langle K_{n}|{\cal L}\varrho|K_{n}\rangle=-i(-1)^{n}\dot{\phi}_{n}(t)\phi_{n}(t),\qquad\qquad\langle K_{n}|\varrho{\cal L}|K_{n}\rangle=+i(-1)^{n}\phi_{n}(t)\dot{\phi}_{n}(t),$ (S8)
[…]
${\cal L}_{cl}=\sum_{n=0}^{D_{\cal O}-1}\frac{\langle K_{n}|\frac{1}{2}\{{\cal L},\varrho\}|K_{n}\rangle}{\langle K_{n}|\varrho|K_{n}\rangle}\;|K_{n}\rangle\langle K_{n}|\,$, (S9)
vanishes identically. Thus ${\cal L}_{nc}={\cal L}$ and using (7)
$\Delta{\cal L}_{nc}=\sqrt{\langle{\cal L}^{2}\rangle-\langle{\cal L}\rangle^{2}}=b_{1}.$ (S10)
[…]
$(\delta_{{\cal L}}A)^{-2}=\sum_{n=0}^{D_{\cal O}-1}\frac{\langle K_{n}|i\,[{\cal L},\varrho]|K_{n}\rangle^{2}}{\langle K_{n}|\varrho|K_{n}\rangle}=4\sum_{n=0}^{D_{\cal O}-1}\dot{\phi}_{n}^{2}(t).$ (S11)
Appendix B Geodesic and Non-geodesic Operator Trajectories on
Let us consider a single qubit with Hamiltonian
$H=\frac{\omega}{2}\sigma_{z}$ (S12)
and initial operator ${\cal O}=\sigma_{x}$. One finds that the Krylov subspace is two-dimensional, spanned by $\{\sigma_{x},\sigma_{y}\}$. The Lanczos coefficients are $b_{1}=\omega$ and $b_{n}=0$ for $n\geq 2$, and the Schrödinger equations become
$\dot{\phi}_{0}=-\omega\phi_{1},\qquad\dot{\phi}_{1}=\omega\phi_{0}.$ (S13)
With $\phi_{0}^{2}+\phi_{1}^{2}=1$, this describes uniform rotation on the unit circle with angular velocity $\omega$. Form this simple example emerges a geometric picture for operator evolution in which the “operator particle” moves along a great circle in the Krylov space with constant speed $b_{1}=\omega$.
In this example since only $b_{1}$ was non-zero the operator moves along the geodesic. To see how non-geodesic motion occurs let us modified the above Hamiltonian by adding a transverse magnetic filed along the $x$ direction
$H=\frac{\omega}{2}\sigma_{z}+\frac{h}{2}\sigma_{x}\,.$ (S14)
with seed operator $\mathcal{O}(0)=\sigma_{x}$. The orthonormal Krylov basis are
$|K_{0}\rangle=\frac{\sigma_{x}}{\sqrt{2}},\qquad|K_{1}\rangle=\frac{i\sigma_{y}}{\sqrt{2}},\qquad|K_{2}\rangle=\frac{\sigma_{z}}{\sqrt{2}},$ (S15)
with Lanczos coefficients $b_{1}=\omega,\,b_{2}=h$ and the wave functions
$\phi_{0}(t)=\cos(\Omega t)+\frac{h^{2}}{\Omega^{2}}(1-\cos\Omega t),\qquad\phi_{1}(t)=\frac{\omega}{\Omega}\sin(\Omega t),\qquad\phi_{2}(t)=\frac{h\omega}{\Omega^{2}}(1-\cos\Omega t),$ (S16)
where $\Omega=\sqrt{\omega^{2}+h^{2}}$. The Lanczos basis closes after three elements with non-zero $b_{2}$ and hence the trajectory is non-geodesic and the bound (17) is satisfied.
Appendix C Rigorous Krylov Light-Cone Bound from Tridiagonality
We present a self-contained derivation of a Lieb-Robinson-type light-cone bound for operator growth in Krylov space. The derivation relies only on the tridiagonal (nearest-neighbor) structure of the Liouvillian in the Krylov basis and standard operator-norm estimates.
Let $\{|K_{n}\rangle\}_{n\geq 0}$ denote the orthonormal Krylov basis generated from the initial operator $|K_{0}\rangle\equiv|{\cal O}(0))$, such that the Liouvillian superoperator $\mathcal{L}$ acts as
$\mathcal{L}|K_{n}\rangle=b_{n+1}|K_{n+1}\rangle+b_{n}|K_{n-1}\rangle,\qquad b_{0}=0,$ (S17)
with real, non-negative Lanczos coefficients $b_{n}$. The Heisenberg evolution of the operator can be written as
$|{\cal O}(t))=e^{-it\mathcal{L}}|K_{0}\rangle.$ (S18)
Projecting onto the Krylov basis defines the Krylov amplitudes $\phi_{n}(t)$ via
$i^{n}\phi_{n}(t)\equiv\langle K_{n}|{\cal O}(t)),\qquad\phi_{n}(t)\in\mathbb{R}.$ (S19)
Equivalently,
$\phi_{n}(t)=i^{-n}\langle K_{n}|e^{-it\mathcal{L}}|K_{0}\rangle.$ (S20)
Since $|i^{-n}|=1$, this phase plays no role in absolute-value bounds.
Expanding the exponential,
$e^{-it\mathcal{L}}=\sum_{m=0}^{\infty}\frac{(-it)^{m}}{m!}\,\mathcal{L}^{m},$ (S21)
one obtains
$\phi_{n}(t)=i^{-n}\sum_{m=0}^{\infty}\frac{(-it)^{m}}{m!}\langle K_{n}|\mathcal{L}^{m}|K_{0}\rangle.$ (S22)
Since the Lanczos representation of $\mathcal{L}$ is tridiagonal, $\mathcal{L}^{m}|K_{0}\rangle$ has support only on Krylov levels $0\leq n\leq m$. The statement follows by induction. For $m=0$, $\mathcal{L}^{0}|K_{0}\rangle=|K_{0}\rangle$. Assume the claim holds for some given fixed $m$. Acting once more with $\mathcal{L}$,
$\mathcal{L}^{m+1}|K_{0}\rangle=\mathcal{L}\!\left(\sum_{n=0}^{m}c_{n}|K_{n}\rangle\right)=\sum_{n=0}^{m}c_{n}\big(b_{n+1}|K_{n+1}\rangle+b_{n}|K_{n-1}\rangle\big),$ (S23)
which has support only on $0\leq n\leq m+1$. $\square$ As a result,
$\langle K_{n}|\mathcal{L}^{m}|K_{0}\rangle=0\qquad\text{for }m<n,$ (S24)
and (S22) reduces to
$\phi_{n}(t)=i^{-n}\sum_{m=n}^{\infty}\frac{(-it)^{m}}{m!}\langle K_{n}|\mathcal{L}^{m}|K_{0}\rangle.$ (S25)
To bound the matrix elements, we use the operator norm
$\|\mathcal{L}\|\equiv\sup_{\|\psi\|=1}\|\mathcal{L}|\psi\rangle\|.$ (S26)
For normalized states $|\phi\rangle$ and $|\psi\rangle$,
$|\langle\phi|\mathcal{L}^{m}|\psi\rangle|\leq\|\mathcal{L}^{m}\|\leq\|\mathcal{L}\|^{m},$ (S27)
where we used submultiplicativity of the operator norm.
On the other hand form Lanczos recursive relation each application of $\mathcal{L}$ maps $|K_{n}\rangle$ to a linear combination of $|K_{n-1}\rangle$ and $|K_{n+1}\rangle$ with coefficients $b_{n},b_{n+1}$. Hence a simple row-sum estimate (equivalently, Gershgorin’s circle theorem) yields
$\|\mathcal{L}\|\leq v_{\mathrm{op}}\equiv\max_{n}(b_{n}+b_{n+1}).$ (S28)
Consequently,
$|\langle K_{n}|\mathcal{L}^{m}|K_{0}\rangle|\leq(v_{\mathrm{op}})^{m}.$ (S29)
Taking absolute values and inserting the norm bound, one arrives at the following bound
$|\phi_{n}(t)|\leq\sum_{m=n}^{\infty}\frac{t^{m}}{m!}\,|\langle K_{n}|\mathcal{L}^{m}|K_{0}\rangle|\leq\sum_{m=n}^{\infty}\frac{(v_{\mathrm{op}}t)^{m}}{m!}.$ (S30)
Using the standard bound on the tail of the exponential series,
$\sum_{m=n}^{\infty}\frac{x^{m}}{m!}\leq\frac{x^{n}}{n!}e^{x},$ (S31)
one obtains
$|\phi_{n}(t)|\leq\frac{(v_{\mathrm{op}}t)^{n}}{n!}\,e^{v_{\mathrm{op}}t}.$ (S32)
Applying Stirling’s approximation, $n!\sim n^{n}e^{-n}\sqrt{2\pi n}$ for large $n$, gives
$|\phi_{n}(t)|\lesssim\exp\!\left[-n\log\!\left(\frac{n}{v_{\mathrm{op}}t}\right)+n+v_{\mathrm{op}}t\right].$ (S33)
Thus $|\phi_{n}(t)|$ decays exponentially once
$n\gtrsim c\,v_{\mathrm{op}}t,$ (S34)
with $c\simeq 3.6$.
This establishes a Lieb–Robinson–type light cone in Krylov space: for any fixed $t$, the amplitude at Krylov levels $n$ beyond this scale is exponentially suppressed. We note that this bound is similar to that of Jacobi operators bound known is the mathematics literature Teschl (2000).
Appendix D Light–cone Bounds for Models With Exact Krylov Wavefunctions
In this appendix we evaluate the geometric bound of Eq. (22) for two important families of models in which the Lanczos coefficients $\{b_{n}\}$ and the Krylov wavefunctions $\{\phi_{n}(t)\}$ are known exactly. These examples serve two purposes: (i) they confirm the sharpness of the geometric light–cone bound, and (ii) they illustrate explicitly the distinction between the front velocity in index space and the true causal velocity set by the arc–length constraint $\ell_{K}(t)=b_{1}t$.
D.1 Models
Chaotic-type model:
The first solvable family consists of models whose Lanczos coefficients take the exact form Parker et al. (2019); Dymarsky and Gorsky (2020)
$b_{n}\;=\;\alpha\sqrt{n(n-1+\eta)},\qquad\alpha>0,\;\eta>0.$ (S35)
The corresponding Krylov amplitudes are
[…]
where $(\eta)_{n}$ is the Pochhammer symbol.
To determine the typical Krylov index $n_{\rm peak}(t)$, we maximize $p_{n}(t)=|\phi_{n}(t)|^{2}$. Using Stirling asymptotics together with
$\frac{(\eta)_{n}}{n!}\sim\frac{n^{\eta-1}}{\Gamma(\eta)}\qquad(n\gg 1),$ (S37)
we obtain
$\ln p_{n}(t)$ $\approx 2n\ln(\tanh\alpha t)-2\eta\ln\cosh(\alpha t)$
$+(\eta-1)\ln n-\ln\Gamma(\eta).$ (S38)
Differentiating w.r.t. $n$ and setting to zero gives the saddle condition
$n_{\rm peak}(t)\;\approx\;-\frac{\eta-1}{2\ln(\tanh\alpha t)}.$ (S39)
For large times one gets
$n_{\rm peak}(t)\;\approx\;\frac{\eta-1}{4}\,e^{2\alpha t}.$ (S40)
Thus the intensity peak propagates exponentially fast in Krylov index. This is consistent with the heuristic front equation
$\dot{n}_{\rm peak}(t)\sim b_{n_{\rm peak}}\sim\alpha n_{\rm peak},$ (S41)
whose solution is $n_{\rm peak}(t)\propto e^{\alpha t}$ (up to a model dependent factor of $2$). This is the hallmark behaviour of chaotic systems with asymptotically linear Lanczos coefficients $b_{n}\sim\alpha n$.
Oscillator-type model:
The second class of solvable models corresponds to harmonic-oscillator–like Lanczos growth $b_{n}\sim\alpha\sqrt{n}$. In these cases one finds Caputa et al. (2022)
$\phi_{n}(t)=\exp\!\left[-\frac{(\alpha t)^{2}}{2}\right]\,\frac{(\alpha t)^{n}}{\sqrt{n!}},\qquad b_{n}=\alpha\sqrt{n}.$ (S42)
This is precisely the coherent-state expansion in the number basis. It yields a Poisson distribution with mean
$n_{\rm peak}(t)=(\alpha t)^{2}.$ (S43)
Thus the front propagates quadratically in time, characteristic of models with $b_{n}\sim\sqrt{n}$.
D.2 Geometric arc-length bound
Linear growth: $b_{n}\approx\alpha n$.
For large $n$,
$\sum_{m=1}^{n(t)}\frac{1}{b_{m}}\sim\frac{1}{\alpha}\sum_{m=1}^{n(t)}\frac{1}{m}\sim\frac{1}{\alpha}\big(\ln n(t)+\gamma\big).$ (S44)
From (22) we obtain
$\frac{1}{\alpha}\ln n(t)\lesssim t\quad\Rightarrow\quad n(t)\;\lesssim\;e^{\alpha t},$ (S45)
i.e. exponential growth of the front which matches the exact behaviour of the Meixner solution (S40), confirming that the geometric inequality is sharp (up to $O(1)$ prefactors independent of $t$). It is worth also noting that linear growth can also been seen in many free theories too Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Vasli et al. (2024).
Square-root growth: $b_{n}\approx\alpha\sqrt{n}$.
Using a Riemann sum,
$\sum_{m=1}^{n(t)}\frac{1}{b_{m}}\sim\frac{1}{\alpha}\sum_{m=1}^{n(t)}m^{-1/2}\sim\frac{2}{\alpha}\sqrt{n(t)}\,$, (S46)
and from (22) we get
$\frac{2}{\alpha}\sqrt{n(t)}\lesssim t\quad\Rightarrow\quad n(t)\;\lesssim\;\frac{(\alpha t)^{2}}{4},$ (S47)
which up to the prefactor $1/4$, this agrees with the exact peak motion $n_{\rm peak}(t)=(\alpha t)^{2}$ for coherent-state models.
It is worth emphasizing that the estimate $\Delta t_{n}\sim 1/b_{n}$ is parametric so that unknown dimensionless prefactors (of order unity) are omitted. For rigorous inequalities one must track these prefactors or derive bounds on the matrix elements $\langle K_{n+1}|\mathcal{L}|K_{n}\rangle=b_{n+1}$. Moreover, the local-time bound (22) is generally sharper than global-operator-norm tail bounds (which involve $v_{\mathrm{op}}=\max_{n}(b_{n}+b_{n+1}))$, because it captures the local structure of the Lanczos chain. Of course both viewpoints are useful, while the geometric identity $\ell_{K}=b_{1}t$ is exact and basis-independent; the local-time sum gives a direct, dynamical prediction for the front $n(t)$ and reproduces known behaviors in soluble models.
The key point illustrated by these examples is the following: Even though the front $n(t)$ may grow faster than linearly in time (quadratically or exponentially), the motion on the Krylov sphere is bounded exactly by a linear geometric arc–length $\ell_{K}=b_{1}t$. Hence the true causal light cone is linear, and its velocity is precisely $v_{\rm op}=b_{1}$.
Large Lanczos coefficients at high $n$ do not endanger causality because those regions of the Krylov graph are dynamically inaccessible until the wave front reaches them; and reaching them requires arc–length, which is generated only at rate $b_{1}$. Thus the geometric light cone remains linear even when the front in index space accelerates.